3.1 mol of fecl3 and 3.2 mol of NH4SCN are added to 1 litre of water. At equilibrium 3.0 mole of FeSCN2+ are formed . The equilibrium constant kc of the reaction Fe3+ +SCN-==FeSCN2+ will be1. 6.66*10-32. 0.303. 3.304. 150

Grade:12th pass

2 Answers

Aayush Parakh
6 years ago
kc = 3  / 3.1*3.2 = 0.30
this is how it is calaculate with help of reaction 
fecl3 dissaciates completely
sidhant
5 years ago
FECL3+NH4SCN GIVES FESCN2+ + 2CL-1 + NH4CL
Reduce other terms you get
Fe3+ + SCN- gives FESCN2+
We know initial mol of fe is 3.1 
initial mol of scn is 3.2
final mol of fescn is 3
final mol of fe =3.1-3=0.1
final mol of scn = 3.2-3 = 0.2
 
now use kc formula
 
Kc= 3/(0.1*0.20 = 150